OpenStudy (anonymous):

How do I integrate ((y^2)*(e^xy))

sam (.sam.):

let u=xy dx=(du)/(y) ye^(u)+C ye^(xy)+C

OpenStudy (anonymous):

i dont get it

sam (.sam.):

Replace xy with u to simplify the integrand. u=xy Since du=y dx, then dx=(du)/((y)). dx=(du)/(y) The indefinite integral of ye^(u) is ye^(u)+C. ye^(u)+C The indefinite integral of y^(2)e^(xy) is ye^(xy)+C. ye^(xy)+C

OpenStudy (anonymous):

thank you a lot!

sam (.sam.):

yw

OpenStudy (dape):

Was the integration with respect to x or to y?

OpenStudy (dape):

.Sam.'s answer is correct if it was with respect to x.

OpenStudy (anonymous):

It is with respect to x, yes :)